# TaskEval: Assessing Difficulty of Code Generation Tasks for Large Language Models

### v2

@@ Abstract.
Large Language Models (LLMs) excel in code-related tasks like code generation, but benchmark evaluations often overlook task characteristics, such as difficulty. Moreover, benchmarks are usually built using tasks described with one single prompt, despite the formulation of prompts having a profound impact on the outcome. This paper introduces a generalist approach, TaskEval, a framework using diverse prompts and Item Response Theory (IRT) to efficiently assess LLMs’ capabilities and benchmark task characteristics, improving the understanding of their performance.
Using two code generation benchmarks, HumanEval+ and ClassEval, as well as 5 code generation LLMs, we show that TaskEval is capable of characterizing the properties of tasks. Using topic analysis, we identify and analyze the tasks of respectively 17 and 21 topics within the benchmarks. We also cross-analyze tasks’ characteristics with programming constructs (e.g., variable assignment, conditions, etc.) used by LLMs, emphasizing some patterns with tasks’ difficulty. Finally, we conduct a comparison between the difficulty assessment of tasks by human-annotators and LLMs. Orthogonal to current benchmarking evaluation efforts, TaskEval can assist researchers and practitioners in fostering better assessments of LLMs. The tasks’ characteristics can be used to identify shortcomings within existing benchmarks. This could be used to generate additional related tasks for the evaluation or improvement of LLM.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Large Language Models (LLMs) have been widely used for various code-related tasks in Software Engineering (SE) (Nguyen and Nadi, 2022; Lemieux et al., 2023; Chen et al., 2023; Moradi Dakhel et al., 2023). Although the performance of LLMs in generating code for different programming languages, such as Python, Java, and C (Yu et al., 2023; Jin et al., 2023; Liu et al., 2023) is promising, the LLM-generated code, like human-written code, is not flawless. Researchers have already acknowledged that LLM-generated code is prone to bugs (Moradi Dakhel et al., 2023; Tambon et al., 2024), and may suffer from vulnerability (Tóth et al., 2024; Majdinasab et al., 2023; Klemmer et al., 2024) and quality issues (Yetiştiren et al., 2023).
@@
Generally, to assess LLMs on code generation tasks and compare them to each other, well-known benchmarks such as HumanEval (Chen et al., 2021; Liu et al., 2023), MBPP (Austin et al., 2021), CoderEval (Yu et al., 2023), and ClassEval (Du et al., 2023) are used. These benchmarks are obtained by handcrafting programming tasks or mining existing public repositories on GitHub and filtering them. In this way, the code generation capability of LLMs in programming tasks is assessed indirectly at the benchmark level based on general characteristics such as granularity level (e.g., function level being less difficult than class level) and through general metrics such as accuracy or Pass@k over the benchmark (Chen et al., 2021). However, this limits the evaluation of the LLMs at the task level as the metric can only report aggregated results over a set of tasks. Moreover, most benchmarks only contain a single prompt for each task within the benchmark and recent studies (Mizrahi et al., 2024; Wang et al., 2022b) showed that the way prompts are formulated has a high impact on the quality of LLM’s output. As such, there is a need for a more fine-grained analysis of the evaluation of the capability of the LLMs at the task-level, that goes beyond general aggregation metrics. Better task-level analysis could also pave the way for more advanced and customized benchmarks, as traditional benchmarks have several limitations such as data contamination, adequate coverage of LLMs’ ability, reliability, and robustness (Jain et al., 2024a; McIntosh et al., 2024; Laskar et al., 2024).
@@
In the field of test theory, Item Response Theory (IRT) (Baker, 2001) was developed precisely to overcome the limitation of considering only an aggregation of the score rather than individual questions. To achieve this, an IRT model estimates a specific ability for each respondent (e.g., a student taking a test or an LLM being evaluated) based on their responses to a set of items (such as questions). These items have in turn some characteristics, for instance, a difficulty level, estimated by the respondents’ answers as well. Contrary to using plain metrics such as Accuracy which do not account for the difference between items (i.e. all LLMs replying correctly on an item vs one LLM replying correctly to it is counted the same for the LLM being correct), IRT will consider each item independently and in conjunction allowing for a more fine-grained analysis. Moreover, by using IRT and the LLMs themselves, we can go beyond using human assessments as a proxy for concepts such as difficulty. Indeed, given that LLMs’ perception of tasks might differ from humans (Ouyang et al., 2024; Kou et al., 2023), LLMs’ grasping of concepts is likely not to be aligned with humans. Recent studies (Vania et al., 2021a; Byrd and Srivastava, 2022; Zhuang et al., 2023) used IRT to assess the characteristics of the questions in a benchmark for an LLM. However, these studies present several limitations. First, they did not focus on code-based benchmarks, which have some differences compared to more classical Natural Language Processing (NLP) tasks. In addition, the assessment is generally conducted via a single prompt, which could affect the output of the LLM. Finally, the studies focused mainly on questions for which a single binary (true/false) answer is given. This limits the application of the methods when multiple code generations should be considered, as is the case when using non-deterministic temperature-based sampling for LLMs. Thus, current benchmarks’ usage and assessment paint a general, yet incomplete picture, as they fail to provide a fine-grained analysis of individual tasks within a benchmark, accounting for variability in the prompt. To address this issue, we propose TaskEval, a framework for the evaluation, identification, and analysis of the difficulty of programming tasks for LLMs.
@@
TaskEval is orthogonal to the usage of existing benchmarks and works on top of available benchmarks. TaskEval starts by generating a variety of prompts for each task on two well-known code generation benchmarks, HumanEval+ (Chen et al., 2021) and ClassEval (Du et al., 2023), representing multiple possible formulations per task. For each programming task, we craft prompts with different levels of context information and distinctive phrasing to represent diverse possible wording of the task using GPT-4, resulting in 18 different prompts per task. The framework then queries an array of 5 LLMs, named CodeLLMs (i.e., CodeLLama 7B (Roziere et al., 2023), MagiCoder 6.7B (Wei et al., 2023), DeepSeekCode 7B (Guo et al., 2024), CodeGemma 7B (Team et al., 2024) and GPT-3.5 (GPT, 2023a)) using the formulated prompts to obtain different outputs for every single task. Those outputs are then used to compute a task score per CodeLLM based on collected metrics from the obtained code snippets. The task scores are then processed using an IRT model to estimate the characteristics of each task (namely difficulty and discriminant) and the abilities of each model. From there, a more nuanced analysis can take place. We highlight the difference between the two benchmarks under study, showing that ClassEval tends to have more extreme tasks in terms of difficulty than HumanEval+. We further the analysis by extracting from the available tasks, respectively 17 and 21 topics for both HumanEval+ and ClassEval, highlighting some topics of interest (such as tasks dealing with sequences of numbers in HumanEval+) that tend to have a higher level of difficulty and be discriminant between CodeLLMs. We also analyze the prevalence of specific program constructs (e.g., variable assignments) by leveraging the Abstract Syntax Tree (AST) of the generated code, uncovering trends related to task characteristics. For example, we observe that the use of conditional statements increases with task difficulty, regardless of code length. Finally, we contrast the difficulty evaluated using TaskEval with human annotators and find a difference between both evaluations, emphasizing that the difficulty of the tasks for CodeLLMs should not be assessed by humans as is. To guide our study, we formulate the following Research Questions (RQs):
@@
The rest of this paper is organized as follows. In Section 3, we describe the methodology we followed to propose TaskEval. Our experimental setup for running TaskEval is detailed in Section 4. We report and then analyze the results obtained in Section 5. A discussion of our results and the insights for the community is presented in Section 6. We conclude the paper in Section 9 after discussing threats to the validity of our findings in Section 8.
@@ 2. Motivating examples
To motivate TaskEval, we will use two examples of tasks. In our study, we make the distinction between task and prompt: the task is the essence of the functionality to be implemented. For instance, “Quicksort” or “Fibonacci sequence” are examples of such tasks. A prompt is a way to formulate the task so it can be understood by and fed to an LLM. For instance, “Write a function to calculate the n-th element of the Fibonacci sequence.” or “Create a function to return the n-th element of the Fibonacci sequence. The Fibonacci sequence is defined as …” are two instances of prompts for the task “Fibonacci sequence”. As such, each task can be expressed by many different (even infinite) prompts, including different wordings, amount of contextual information about the task, like examples, etc. To compare the effect between prompt and task, we will thus calculate the difficulty using our proposed approach (Section 3) and a more traditional approach consisting of an ensemble method (i.e., how many CodeLLMs failed/managed the task). To emphasize the importance of using multiple prompts for estimating the task’s characteristics, our approach employs multiple prompts, in contrast to the traditional method, which relies solely on a single prompt (i.e., the original benchmark prompt). Moreover, to further contrast our results, we will compare with the difficulty as reported by human developers using a crowd-based approach (see Section 5.4).
1def select_by_age_range(self, min_age, max_age):
2 """Generates a SQL statement to select records within a specified age range.
3 :param min_age: int. The minimum age.
4 :param max_age: int. The maximum age.
5 :return: str. The generated SQL statement.
6 $>$ sql.select_by_age_range(20, 30)
7 ’SELECT * FROM table1 WHERE age BETWEEN 20 AND 30;’
8 """
9
10def select_by_age_range(self, min_age, max_age):
11 """ Generate a SQL statement to select records within a specified age range using
12 the "min_age" and "max_age" parameters. Return the generated SQL statement as a string.
13 :param min_age: int. The minimum age.
14 :param max_age: int. The maximum age.
15 :return: str. The generated SQL statement.
16 """
Listing 1: Example of discrepancy on a SQL task. (Top) Original Prompt, (Bottom) One of our generated prompt
@@
In the first example (see Listing 1), which involves generating an SQL statement, all CodeLLMs manage to generate a correct code for this task using the original prompt. However, TaskEval returns a difficulty of $0.75$ (out of $1$) on this task, making it a quite complex task. As one can see from the example prompts, the main difference stems from the fact that our prompts will not contain the example included in the original prompt which basically gives what the SQL statement should be like. As such, without this help, the task becomes much harder for the CodeLLMs. And so, while the original prompt might be considered easy, the task itself is not easy, as many potentially generated prompts would not lead to a correct code. If we compare with human assessment, developers agreement settles on a difficulty of 3/6 (given the scale used in our experiment), meaning it is a task of relative medium difficulty. In the second example (see Listing 2), which involves checking if an array is correctly sorted in ascending order without more than one duplicate, all CodeLLMs failed to generate a correct code when using the single prompt. For all CodeLLMs, the main source of error is the handling of duplicates. We believe this issue arises because CodeLLMs struggle with cases involving “multiple duplicates”, and the provided examples are insufficient. However, by varying the prompts, the difficulty level computed by our approach is $0.17$ — a relatively low value, among the lowest in this benchmark. Moreover, slight modifications to the prompt can lead to better-generated code compared to relying solely on the original prompt. For instance, the example we give in Listing 2 shows that by wording the condition with “appears more than twice”, the CodeLLMs manage to generate more correct codes. Note that, in that case, having an example did not appear to have helped the CodeLLMs generate a correct code, contrary to reformulating or adding information to the prompt. Through these examples, we see that traditional assessment may be limited in accurately evaluating the characteristics of tasks, which can, in turn, hinder benchmark analysis. If we compare with human assessment, developers agreement settles on a difficulty of 4/6 (given the scale used in our experiment), meaning it is also a task of relative medium difficulty. All in all, we observe that difficulty can drastically change for LLMs based on the formulation of the prompts and that it also contrasts with human judgment of the tasks.
@@ 4 Examples
11 """ Write a function named ’is_sorted’ that checks if a given list of non-negative integers is sorted in ascending order. Additionally, the function should return False if any integer appears more than twice in the list.
@@
Listing 2: Example of discrepancy on a Sorting Task. (Top) Original Prompt, (Bottom) One of our generated prompt

### v3

@@ Abstract.
Large Language Models (LLMs) excel in code-related tasks like code generation, but benchmark evaluations often overlook task characteristics, such as difficulty. Moreover, benchmarks are usually built using tasks described with a single prompt, despite the formulation of prompts having a profound impact on the outcome. This paper introduces a generalist approach, TaskEval, a framework using diverse prompts and Item Response Theory (IRT) to efficiently assess LLMs’ capabilities and benchmark task characteristics, improving the understanding of their performance.
Using two code generation benchmarks, HumanEval+ and ClassEval, as well as 8 code generation LLMs, we show that TaskEval is capable of characterising the properties of tasks. Using topic analysis, we identify and analyse the tasks of 17 and 21 topics within the benchmarks. We also cross-analyse tasks’ characteristics with programming constructs (e.g., variable assignment, conditions, etc.) used by LLMs, emphasising some patterns with tasks’ difficulty. Finally, we conduct a comparison between the difficulty assessment of tasks by human annotators and LLMs. Orthogonal to current benchmarking evaluation efforts, TaskEval can assist researchers and practitioners in fostering better assessments of LLMs. The tasks’ characteristics can be used to identify shortcomings within existing benchmarks or improve the evaluation of LLMs.
Large Language Models, Code Benchmark Assessments, Item Response Theory
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Large Language Models (LLMs) have been widely used for various code-related tasks in Software Engineering (SE) (Nguyen and Nadi, 2022; Lemieux et al., 2023; Chen et al., 2023; Moradi Dakhel et al., 2023). Although the performance of LLMs in generating code for different programming languages, such as Python, Java, and C (Yu et al., 2023; Jin et al., 2023; Liu et al., 2023) is promising, the LLM-generated code, like human-written code, is not flawless. Researchers have already acknowledged that LLM-generated code is prone to bugs (Moradi Dakhel et al., 2023; Tambon et al., 2024), and may suffer from vulnerability (Tóth et al., 2024; Majdinasab et al., 2023; Klemmer et al., 2024) and quality issues (Yetiştiren et al., 2023). Generally, to assess LLMs on code generation tasks and compare them to each other, well-known benchmarks such as HumanEval (Chen et al., 2021; Liu et al., 2023), MBPP (Austin et al., 2021), or ClassEval (Du et al., 2023) are used. These benchmarks are obtained by handcrafting programming tasks or mining existing public repositories on GitHub and filtering them. In this way, the code generation capability of LLMs in programming tasks is assessed indirectly at the benchmark level based on general characteristics such as granularity level (e.g., function level being less difficult than class level) and through general metrics such as accuracy or Pass@k over the benchmark (Chen et al., 2021). However, this limits the evaluation of the LLMs at the task level as the metric can only report aggregated results over a set of tasks. Moreover, most benchmarks only contain a single prompt for each task within the benchmark and recent studies (Mizrahi et al., 2024; Wang et al., 2022b) showed that the way prompts are formulated has a high impact on the quality of LLM’s output. As such, there is a need for a more fine-grained analysis of the evaluation of the capability of the LLMs at the task-level, that goes beyond general aggregation metrics. Better task-level analysis could also pave the way for more advanced and customised benchmarks, as traditional benchmarks have several limitations, such as data contamination, inadequate coverage of LLMs’ ability, reliability, and robustness (Jain et al., 2024; McIntosh et al., 2024; Laskar et al., 2024).
@@
In the field of test theory, Item Response Theory (IRT) (Baker, 2001) was developed precisely to overcome the limitation of considering only an aggregation of the score rather than individual questions. To achieve this, an IRT model estimates a specific ability for each respondent (e.g., a student taking a test or an LLM being evaluated) based on their responses to a set of items (such as questions). These items have in turn some characteristics, for instance, a difficulty level, estimated by the respondents’ answers as well. Contrary to using plain metrics such as Accuracy, which do not account for the difference between items (i.e. all LLMs replying correctly on an item vs one LLM replying correctly to it is counted the same for the LLM being correct), IRT will consider each item independently and in conjunction, allowing for a more fine-grained analysis. Moreover, by using IRT and the LLMs themselves, we can go beyond using human assessments as a proxy for concepts such as difficulty. Indeed, given that LLMs’ perception of tasks might differ from humans (Ouyang et al., 2024; Kou et al., 2023), LLMs’ grasping of concepts is likely not to be aligned with humans. Recent studies (Vania et al., 2021a; Byrd and Srivastava, 2022; Zhuang et al., 2023) used IRT to assess the characteristics of the questions in a benchmark for an LLM. However, these studies present several limitations. First, they did not focus on code-based benchmarks, which have some differences compared to more classical Natural Language Processing (NLP) tasks. In addition, the assessment is generally conducted via a single prompt, which could affect the output of the LLM. Finally, the studies focused mainly on questions for which a single binary (true/false) answer is given. This limits the application of the methods when multiple code generations should be considered, as is the case when using non-deterministic temperature-based sampling for LLMs. Thus, current benchmarks’ usage and assessment paint a general, yet incomplete picture, as they fail to provide a fine-grained analysis of individual tasks within a benchmark, accounting for variability in the prompt. To address this issue, we propose TaskEval, a framework for the evaluation, identification, and analysis of the difficulty of programming tasks for LLMs.
@@
TaskEval is orthogonal to the usage of existing benchmarks and works on top of available benchmarks. TaskEval starts by generating a variety of prompts for each task on two well-known code generation benchmarks, HumanEval+ (Chen et al., 2021) and ClassEval (Du et al., 2023), representing multiple possible formulations per task. For each programming task, we craft prompts with different levels of context information and distinctive phrasing to represent diverse possible wording of the task using GPT-4, resulting in 18 different prompts per task. The framework then queries an array of 8 LLMs of varying size, architecture and training process, named CodeLLMs (i.e., YiCoder-1.5B (AI et al., 2025), CodeLLama 7B (Roziere et al., 2023), MagiCoder 6.7B (Wei et al., 2023), DeepSeekCoder 7B (Guo et al., 2024), CodeGemma 7B (Team et al., 2024), QwenCoder-2.5 7B (Yang et al., 2024), GPT-3.5 (GPT, 2023a), DeepSeek-V3 (DeepSeek-AI, 2024)) using the formulated prompts to obtain different outputs for every single task. Those outputs are then used to compute a task score per CodeLLM based on collected metrics from the obtained code snippets. The task scores are then processed using an IRT model to estimate the characteristics of each task (namely, difficulty and discriminant) and the abilities of each model. From there, a more nuanced analysis can take place. We highlight the difference between the two benchmarks under study, showing that ClassEval tends to have more extreme tasks in terms of difficulty (easy and difficult) than HumanEval+. We further the analysis by extracting from the available tasks, respectively 17 and 21 topics for both HumanEval+ and ClassEval, highlighting some topics of interest (such as tasks dealing with sequences of numbers in HumanEval+) that tend to have a higher level of difficulty and be discriminant between CodeLLMs. We also analyse the prevalence of specific program constructs (e.g., variable assignments) by leveraging the Abstract Syntax Tree (AST) of the generated code, uncovering trends related to task characteristics. For example, we observe that the use of conditional statements increases with task difficulty, regardless of code length. Finally, we contrast the difficulty evaluated using TaskEval with human annotators and find a difference between both evaluations, emphasising that the difficulty of the tasks for CodeLLMs should not be assessed by humans as is. To guide our study, we formulate the following Research Questions (RQs):
@@
The rest of this paper is organised as follows. In Section 3, we describe the methodology we followed to propose TaskEval. Our experimental setup for running TaskEval is detailed in Section 4. We report and then analyse the results obtained in Section 5. A discussion of our results and the insights for the community is presented in Section 6. We conclude the paper in Section 9 after discussing threats to the validity of our findings in Section 8.
@@ 2. Motivating examples
To motivate TaskEval, we will use two examples of tasks. In our study, we make the distinction between task and prompt: the task is the essence of the functionality to be implemented. For instance, “Quicksort” or “Fibonacci sequence” are examples of such tasks. A prompt is a way to formulate the task so it can be understood by and fed to an LLM. For instance, “Write a function to calculate the n-th element of the Fibonacci sequence.” or “Create a function to return the n-th element of the Fibonacci sequence. The Fibonacci sequence is defined as …” are two instances of prompts for the task “Fibonacci sequence”. As such, each task can be expressed by many different (even infinite) prompts, including different wordings, amount of contextual information about the task, like examples, chain-of-thoughts, etc. To emphasise the importance of using multiple prompts for estimating the task’s characteristics, our approach employs multiple prompts, in contrast to the traditional method, which relies solely on a single prompt.
@@
In the first example (see Listing 1), extracted from HumanEval+, which involves checking if an array is correctly sorted in ascending order without more than one duplicate, most CodeLLMs failed to generate a correct code when using the original prompt. Pass@1 on this prompt is 0 for all models except DeepSeek-V3 with 1.0 and QwenCoder with 0.2. The main source of error is the handling of duplicates. We believe this issue arises because most CodeLLMs struggle with cases involving “multiple duplicates”, and the provided examples are insufficient. However, the difficulty level computed by our approach is $0.17$ — a relatively low value, among the lowest in this benchmark. Indeed, by slightly modifying the prompt, better-generated code can be obtained compared to the one obtained with the original prompt. For instance, the example we give in Listing 1 shows that by wording the condition with “occurs more than twice”, the CodeLLMs manage to generate more correct codes. In that case, the Pass@1 improves drastically, ranging from 0.6 to 1.0 for all models. Note that, in that case, having an example did not appear to have helped the CodeLLMs generate a correct code, contrary to reformulating the prompt.
@@ 4 Examples
11 """ Create a function called ’is_sorted’ to determine if a list of non-negative integers is in ascending order. The function should also return False if any integer in the list occurs more than twice.
@@
Listing 1: Example of discrepancy on a Sorting Task. (Top) Original Prompt, (Bottom) One of our generated prompt.
In the second example (see Listing 2), extracted from ClassEval, which involves coding a part of a basic restaurant management system with the method ‘add_dish‘, we generated two prompts starting from an original one: the first one contains relatively few information (top), obtained by having the prompt simply reformulate the original prompt; the second one (bottom) contains more information by adding natural description of the inner working of the function, obtained by describing partly the available oracle. Both prompts also included the class constructor and description to provide the CodeLLMs with general context. In the first case, the Pass@1 was 1.0 only for DeepSeek-V3 and QwenCoder, the rest of the models oscillating between 0 and 0.2. Several CodeLLMs struggled on multiple points: the implicit fact that, if the dish is added, the quantity should be lowered, hallucinating undefined methods, handling the non-required sales, etc. However, our approach evaluates the difficulty of the approach to $0.2$, relatively low difficulty. Indeed, if instead a more thorough, yet high-level, description of the intended functionality is given, such as is the case with the bottom prompt, the Pass@k improves for all models, reaching between $0.6$ and $1.0$.
1
2class Order:
3 """
4 The class manages restaurant orders by allowing the addition of dishes, calculation of the total cost, and checkout.
5 """
6
7 def __init__(self):
8 """
9 Initialize the order management system
10 self.menu stores the dishes of resturant inventory
11 menu = [{"dish": dish name, "price": price, "count": count}, ...]
12 self.selected_dishes stores the dished selected by customer
13 selected_dish = {"dish": dish name, "count": count, price: price}
14 self.sales stores the sales of each dish
15 sales = {dish name: sales}
16 """
17 self.menu = []
18 self.selected_dishes = []
19 self.sales = {}
20
21 ...
22
23 # Function to implement
24 def add_dish(self, dish):
25 """ Append the dish to ’self.selected_dishes’ provided that the dish is present in ’self.menu’ and adequate quantities are available. Return ’True’ on successful addition, and ’False’ otherwise.
26 :param dish: dict, the information of dish. dish = {"dish": dish name, "count": count, price: price}
27 :return: True if successfully added, or False otherwise.
28 """
29
30 def add_dish(self, dish):
31 """ Incorporate the dish into ’self.selected_dishes’ only if it is present in ’self.menu’ with sufficient quantities. Examine each dish in ’self.menu’ to see if the name matches the provided dish. If a match is discovered and the current quantity exceeds or meets the needed quantity, update the stock by decreasing the quantity in ’self.menu’, then add it to ’self.selected_dishes’. If the addition is successful, return ’True’; otherwise, return ’False’.
32 :param dish: dict, the information of dish. dish = {"dish": dish name, "count": count, price: price}
33 :return: True if successfully added, or False otherwise.
34 """
Listing 2: Example of discrepancy on a SQL task. For ”Function to implement”: (Top) Low-level information prompt generated, (Bottom) Higher-level information prompt generated
Through these examples, we see that traditional assessment may be limited in accurately evaluating the characteristics of tasks, which can, in turn, hinder benchmark analysis. More generally, if we were to compute the pass@1 per task for different rephrasing, while we would observe limited change in the average Pass@1 (0.01 to 0.03 standard deviation for HumanEval+ across models and rephrasing and 0.02 to 0.05 for ClassEval), the Pass@1 on the same task can double from 10% to 30% of the benchmarks tasks for a given model, that is the correctness can be drastically altered by the prompt formulation. We observe similar results if we consider the level of information. In those cases, it is then not possible to infer the “difficulty” of the task; at best, the difficulty of a given prompt formulation could be deduced. Confronted with this issue, our approach aims to account for this disparity by giving a difficulty score that accounts for possible prompt variation as well as disparity between models.